OpenStudy (anonymous):

Question 4 Evaluate f(n) = -3n3 + 5n2 + n + 6 at n = 2 A 82 B 52 C Simplify (9 + 7i)/(-10 + 8i)

OpenStudy (anonymous):

\[f(n) = -3n^{3} +5n^{2} +n+6 \] To solve for f(n=2), you simply have to plug the number 2 in for every n in the equation: \[f(n) = -3(2)^{3} +5(2)^{2} +(2)+6 \] Now you just simplify and add up the terms. \[f(n) = -3*8 +5*4 +(2)+6 \] \[f(n) = -24 +20 +(2)+6 \] I'm sure you can find the answer from there.

OpenStudy (anonymous):

For the complex number, \[(9+7i)(-10+8i)\] you solve this like you would any algebra equation and treat i as a variable at the beginning (spell check?) \[9*(-10)+9*(8i)+7i(-10)+7i(8i)\] Simplify \[90+72i+-70i+56i^{2}\] now you combine like terms \[90+2i+56i^{2}\] now the complex number comes in, remember that \[\sqrt{-1}=i\] therefore \[i^{2}=(\sqrt{-1})^{2}=-1\] If you plug that back in for \[i^{2}\] then \[90+2i+56(-1)\] simplify once more and you get your answer \[34+2i\]